OpenStudy (anonymous):

y = ax^2 + bx + c Determine the solution set of 0 = ax^2 + bx + c. is it {-2} ?

mathslover (mathslover):

do you know how to calculate the solution set?

OpenStudy (anonymous):

i yesterday posted solution o the same problem. you can check it here http://openstudy.com/study#/updates/5015fdf9e4b0fa246733cc56

Parth (parthkohli):

It is actually a lot more related to the quadratic formula.

Parth (parthkohli):

In a quadratic equation \(ax^2 + bx + c = 0\), there are two roots. Do you know the quadratic formula?

OpenStudy (anonymous):

@helpme9283 did you get it?

Parth (parthkohli):

Remember: the two roots are given as follows. \[\alpha = {-b + \sqrt{b^2 - 4ac} \over 2a} \] \[\beta = {-b - \sqrt{b^2 - 4ac}\over 2a} \]

Parth (parthkohli):

Does that give you a nice hint?

OpenStudy (anonymous):

here it is when y=0 ax^2+bx+c=0 then using by completing square method solution is as follows.

Parth (parthkohli):

Solution set consists of numbers that are solutions. For example, if 2 and 3 are the solutions, then the solution set it \((2,3)\).

Parth (parthkohli):

What are the solutions that I gave you?

OpenStudy (anonymous):

2&3

Parth (parthkohli):

No. What are the solutions that I gave you for \(ax^2 + bx + c = 0\)?

OpenStudy (anonymous):

i didnt get it..

Parth (parthkohli):

Read the post where I mentioned alpha and beta.

OpenStudy (anonymous):

this is the solution set. \[\Large [{\frac{-b+\sqrt{b^2-4ac}}{2a}} ,{\frac{-b-\sqrt{b^2-4ac}}{2a}}]\]

Parth (parthkohli):

Oh-oh.

mathslover (mathslover):

or : \[\LARGE{roots=\frac{-b\pm \sqrt{b^2-4ac}}{2a}}\]

Parth (parthkohli):

But that is not the solution set.

Parth (parthkohli):

@sami-21 Is right, but should consider not posting answers :)